# Wall-resolved large eddy simulations of a pitching NACA0012 incurring in deep dynamic stall

### v1

@@ Abstract
This study investigates the flow evolution past a pitching NACA0012 airfoil undergoing deep dynamic stall using a wall-resolved large eddy simulation approach. Numerical results are validated against experimental data from Lee and Gerontakos (2004) at Reynolds number Re = 135 000 and a reduced frequency $k$ = 0.1. A comprehensive analysis of the computational grid span size is presented, highlighting the critical requirement for a span-to-chord ratio of at least one. Furthermore, a comparative assessment with state-of-the-art RANS and hybrid RANS/LES turbulence models is conducted to identify the sources of potential discrepancies between the numerical simulations and the experimental data. All the numerical models demonstrate the same behavior and exhibit similar differences with the experiments.
1 Introduction
@@
Dynamic stall presents a significant challenge within the domain of aerodynamic design. It manifests as a transient, unsteady phenomenon impacting airfoils subjected to rapid variations in angle of attack. This behavior deviates considerably from the well-established theory of static stall, where a progressive reduction in lift occurs with increasing angle of attack. The pivotal factor in dynamic stall is the disruption of the attached boundary layer. This triggers the formation of the dynamic stall vortex (DSV) that convects downstream along the suction surface of the airfoil. The subsequent development and interaction of DSV with the boundary layer lead to complex flow separation phenomena (Marilyn J. et al., 2020). Dynamic stall induces flow separation over a substantial portion of the airfoil at angle of attacks exceeding the critical static stall limit. The consequence of this extensive flow separation is a drastic, and often delayed, reduction in lift, potentially accompanied by a significant increase in drag and pitch down moment. This behavior can have a profound impact on the performance and stability of aerodynamic structures. Notably, dynamic stall plays a crucial role in several key areas such as: helicopter rotor blades, wind turbines, and high-performance aircraft maneuvers.
Figure 1: Sketch of dynamic stall over a pitching airfoil
@@
Focusing on the physics of the dynamic stall problem in pitching airfoils, it is possible to identify five main phases. Figure 1 presents a sketch of the flow in each of these stages. During the initial portion of the cycle, the airfoil experiences attached flow due to the low angles of incidence. As the angle of attack increases throughout the upstroke phase, a laminar separation bubble forms on the profile. Concurrently, the boundary layer transitions from laminar to turbulent flow, initiating at the trailing edge and progressing forward. The interaction between these two phenomena triggers the formation of the dynamic stall vortex, which subsequently convects rearward. Upon reaching the maximum angle of incidence and at the beginning of the downstroke phase, the flow over the airfoil suction side becomes fully stalled. Reattachment starts at a reduced angles of attack due to the negative pitch rate which increases the effective angle of attack.
@@
The complex and transient nature of dynamic stall presents a significant obstacle in computational modeling and design optimization (Gardner et al., 2023; Baldan et al., 2024). Research efforts are primarily focused on developing robust high-fidelity computational fluid dynamics (CFD) tools and low-fidelity reduced order models (ROMs) to be used in flow control strategies. Indeed, by gaining a deeper understanding of dynamic stall and its impact, the design and control of aerodynamic structures can be improved (Gardner et al., 2019; Hariharan et al., 2014). This will ultimately lead to safer and more efficient flight, improved wind energy generation, and robust performance across diverse operating conditions.
Looking at high-fidelity simulations, many contributions using large eddy simulations (LES) can be found in the literature. In particular, several aspects have been numerically investigated for finite wings ranging from aspect ratio (Hammer et al., 2022) to sweep angle (Hammer et al., 2023) and compressibility effects (Benton and Visbal, 2020). Another aspect of dynamic stall study is the analysis of nominal infinite wings obtained imposing periodic boundary conditions in the spanwise direction. According to the best practice for detached-eddy simulations (DES), presented in Shur et al. (1999), to reproduce the DSV is necessary a spanwise size of at least one chord. Indeed, it is well known that the characteristic dimension of the DSV is on the order of the airfoil chord length when incurring in deep dynamic stall (McCroskey, 1981). Despite the just mentioned considerations, a significant portion of LES literature employs extruded airfoil geometries with spanwise lengths ranging from only 5% to 20% of the chord size.
In Guillaud et al. (2018), a pitching NACA0012 airfoil at Re $=2\cdot 10^{4}$ is investigated using a numerical model with spanwise length of $0.5c$. In particular, the analysis focuses into the influence of the reduced frequency on the aerodynamic coefficients drop delay in comparison with the static case.
Other works on dynamic stall phenomenology and control have been developed by Benton and Visbal (2019b, a); Visbal and Benton (2018). They leverage high-frequency control to mitigate the strong burst of the aerodynamic loads due to the leading edge vortex (LEV) detachment. The effect of the Reynolds number on the laminar separation bubble (LSB) is also part of the investigation. However, the airfoil extrusion in all the contributions is limited to $0.05c$, $0.1c$ and $0.2c$.
Recently, Lee and Gross (2024) tried to match other LES results present in the literature reproducing a ramp-up motion with spanwise size equal to $0.1c$ and a sinusoidal pitching test case using a $0.02c$ wide grid. The former is in good agreement with the reference data of Gupta and Ansell (2019) while the latter shows large discrepancies in the downstroke phase compared to Benton and Visbal (2019b).
On the contrary, Visbal and Garmann (2018) analyze the span effect starting from $0.1c$ and up to $1.6c$ in a ramp-up motion. They showed that there is an effect when the DSV approaches the trailing edge of the profile and starts to interact with the trailing edge vortex (TEV). A difference is present also in the loads especially for the moment coefficient. Despite the span analysis at the beginning of the work, they focused their investigation on the LSB and DSV generation and all the following displayed results refer to the $0.1c$ case.
Another contribution by the same authors, Visbal and Garmann (2019), focuses on the numerical simulations of typical experimental setups. Namely, the aspect ratio of the wing section together with end walls and end tips is considered. They provide a quantification of the discrepancy magnitude among all the configurations, but they do not report any comparison with real experimental data.
Less recent contributions tried to match experiments with large eddy simulations. Patil et al. (2018) delves into the analysis of an H-Darrieus rotor trying to match wind tunnel data. They were able to capture the attached flow behavior but not the stalled region. The span size is limited to $0.05c$.
Other examples are Badoe et al. (2019); Kasibhotla and Tafti (2014) that investigates pitching NACA0012 airfoils at transitional Reynolds number using meshes with spanwise length of $0.25c$ and $0.18c$, respectively.
@@
Another critical aspect in the use of high-fidelity simulations to model dynamic stall phenomena is the difference present between numerics and experiments. If other numerical approaches are considered, such as unsteady Reynolds-Averaged Navier-Stokes (RANS) models and hybrid RANS/LES techniques, an even larger number of works is present in literature (Kim and Xie, 2016; Karbasian and Kim, 2016; Khalifa et al., 2021, 2023; Geng et al., 2018; Zanotti and Gibertini, 2013; Zanotti et al., 2014b, a; Avanzi et al., 2022, 2021; Gardner and Richter, 2020; Baldan and Guardone, 2024a, b). Most of this works conclude that the discrepancies between experimental data and numerical simulations rely on the adopted turbulence model to close the RANS system. A vast number of turbulence models have been adopted ranging from one equation Spalart-Allmaras formulation (Spalart and Allmaras, 1992) to full Reynolds stress model (RSM) (Launder et al., 1975), and including the well-known $k-\epsilon$ (Launder and Spalding, 1974) and $k-\omega$ SST (Menter, 1994). Also, transition models are leveraged, for instance the recent $k-\omega$ SST with intermittency equation (Menter et al., 2004). Regarding hybrid strategies, common choices are detached eddy simulations (DES) (Spalart et al., 1997) and their improved versions (DDES, IDDES). More recent approaches are scale adaptive simulations (SAS) (Menter and Egorov, 2010) and stress-blended eddy simulations (SBES) (Menter, 2018).
@@
This work proposes a numerical investigation of a NACA0012 airfoil undergoing sinusoidal pitching motion at a Reynolds number of 135 000 and a reduced frequency of k = 0.1. The study utilizes a wall-resolved large eddy simulation approach to analyze the influence of spanwise length on the aerodynamic behavior during both the upstroke and downstroke phases. Span sizes ranging from $0.2c$ to $1.2c$ are examined. Furthermore, the obtained numerical results are compared with the experimental data proposed by Lee and Gerontakos (2004). Additionally, a comparison will be presented with current state-of-the-art RANS and hybrid RANS/LES models. This comparative analysis aims to elucidate whether the observed discrepancies between the numerical results and the experimental data stem from the turbulence model employed or other effects.
@@ 2 Numerical setup
Figure 2: Three-dimensional mesh of the wing section with span-to chord ratio of 1.2.
The present numerical simulations replicate the experimental conditions presented in Lee and Gerontakos (2004). A NACA0012 airfoil with a 0.15 m chord underlying a sinusoidal pitching motion defined as:
@@ 2.1 Grid generation
Table 1: Grid specifications. Size of elements is reported in meters.
@@
Span size $\mathbf{N_{\xi}}$ $\mathbf{N_{\eta}}$ $\mathbf{N_{z}}$ Nodes Hexahedra $\mathbf{\Delta\xi_{\text{LE}}}$ $\mathbf{\Delta\xi_{\text{TE}}}$ $\mathbf{\overline{\Delta\xi}_{\text{suction}}}$ $\mathbf{\overline{\Delta\xi}_{\text{pressure}}}$
$0.2c$ 934 334 201   62.7M   62.1M $1.00\cdot 10^{-4}$ $6.52\cdot 10^{-5}$ $2.82\cdot 10^{-4}$ $4.04\cdot 10^{-4}$
$0.4c$ 401 125.1M 124.3M - - - -
$0.8c$ 801 249.9M 248.6M - - - -
$1.2c$ 1201 374.7M 372.8M - - - -
@@
The CFD mesh is generated using Fidelity Pointwise v2023.2 Cadence (2023). To avoid the non-physical flow patterns caused by the typical C-grid strong anisotropic elements, especially near the trailing edge Baldan and Guardone (2024b), the O-grid is adopted in this research. Table 1 reports the details of the grids, where $\xi$ is the coordinate over the profile, $\eta$ is the direction normal to the surface, and $z$ is the spanwise direction.
@@
The O-grid is obtained through a hyperbolic extrusion in the normal direction starting from the profile surface. The wall-normal spacing follows a geometric progression with a growth rate of 1.03. The standard values required for Wall-Resolved Large Eddy Simulation (WRLES), typically demanding resolution values of $\xi^{+}<40$, $\eta^{+}<5$, and $z^{+}<20$ Piomelli and Balaras (2002); Georgiadis et al. (2010), are always satisfied at the wall, even for the most demanding flow condition when the acceleration, due to the high angle of attack, increases the velocity at more than twice the free-stream value. Thus, the first layer is positioned at $1.5\cdot 10^{-5}\ \text{m}$ over the surface. Another critical aspect in the mesh to capture dynamic stall phenomena is the element size at the leading and at the trailing edges. The leading edge spacing is extremely important since it covers the region where the laminar separation bubble resides during the pitch up phase and subsequently the dynamic stall vortex is generated. The same consideration can be applied also to the trailing edge since stall also starts from the back of the profile. Also for this reason the blunt trailing edge is jointed with a circle that avoids the imposition of the separation point and at the same time increases the orthogonality of the grid allowing a faster convergence. The element size at the nose of the airfoil is $\Delta\xi_{\text{leading edge}}=1.00\cdot 10^{-4}\ m$ while at the trailing edge is $\Delta\xi_{\text{trailing edge}}=6.52\cdot 10^{-5}\ m$. In order to reduce the number of points of the meshes the pressure side of the airfoil, which is not directly involved in the DSV advection, is discretized with a lower number of points. Specifically, the suction side has 550 nodes, the pressure side instead is limited to 385. This results in an average spacing on the upper surface of $\overline{\Delta\xi}_{\text{suction}}=2.82\cdot 10^{-4}\ m$ and on the lower surface of $\overline{\Delta\xi}_{\text{pressure}}=4.04\cdot 10^{-4}\ m$. In all the meshes the farfield is placed at $40c$ from the airfoil to reduce the boundary influences.
@@
The three-dimensional grid is obtained through a uniform extrusion of the two-dimensional O-grid in the $z$ direction. The number of points is retrieved respecting the LES requirements (Choi and Moin, 2012). For the smallest mesh, extending for $0.2c$ in span, the number of points is equal to 201. When moving to larger grids, the element size is kept the same resulting in 401, 801, and 1201 nodes for meshes with span size of $0.4c$, $0.8c$, and $1.2c$, respectively. Figure 2 illustrates the three-dimensional grid with $1.2c$ span size.

### v2

# Wall-resolved large eddy simulations of a pitching airfoil incurring in deep dynamic stall

@@ Abstract
This study investigates the flow evolution around a sinusoidal pitching NACA 0012 airfoil, defined by the National Advisory Committee for Aeronautics (NACA), undergoing deep dynamic stall using a wall-resolved large eddy simulation (LES) approach. Numerical results are assessed against experimental data from Lee and Gerontakos (2004) at Reynolds number Re = 135 000 and reduced frequency $k$ = 0.1. A comprehensive analysis of the computational model span size is presented, highlighting the requirement for a span-to-chord ratio of at least one to correctly capture the dynamic stall vortex physics in the downstroke phase. Furthermore, a comparative assessment with state-of-the-art Reynolds-Averaged Navier-Stokes (RANS), hybrid RANS/LES, and the experimental data is carried out. All the numerical models concur to the same flow behavior and exhibit similar differences with the experiments.
@@
Dynamic stall is a transient unsteady phenomenon impacting airfoils subjected to rapid variations in the angle of attack. It deviates significantly from the well-established theory of static stall, where an abrupt reduction in lift occurs once a limit value of the angle of attack, the stall angle, is exceeded. In dynamic stall, the separation of the attached boundary layer triggers the formation of the dynamic stall vortex (DSV) that convects downstream along the suction surface of the airfoil. The subsequent development and interaction of the DSV with the boundary layer leads to complex flow separation phenomena (Marilyn J. et al., 2020). Dynamic stall induces flow separation over a substantial portion of the airfoil at angles of attack exceeding the critical static stall limit. The extensive flow separation results in a drastic, and often delayed, reduction in lift, combined with a significant increase in drag and pitch down moment. Dynamic stall can have a profound impact on the performance and stability of aerodynamic structures. Notably, dynamic stall plays a crucial role in several key areas: helicopter rotor blades, wind turbines, and high-performance aircraft maneuvers. The occurrence of dynamic stall limits the maximum speed of conventional rotorcraft.
Figure 1: Sketch of dynamic stall over a pitching airfoil: a) attached flow at small angles of incidence; b) laminar separation bubble generates increasing the incidence of the profile; c) dynamic stall vortex detaches and convects backward; d) complete stalled flow over the airfoil during the downstroke phase; e) flow reattachment at reduced angle of attack.
@@
With reference to Figure 1, it is possible to identify five main phases of the dynamic stall. a) During the initial portion of the cycle, the airfoil experiences attached flow due to the low angles of incidence. b) As the angle of attack increases throughout the upstroke phase, a laminar separation bubble forms on the airfoil. Concurrently, the boundary layer transitions from laminar to turbulent flow, initiating at the trailing edge and progressing downstream. c) The interaction between the boundary layer and the laminar separation bubble triggers the formation of the dynamic stall vortex, which eventually convects rearward. d) Upon reaching the maximum angle of incidence and at the beginning of the downstroke phase, the flow over the airfoil suction side becomes fully stalled. e) Reattachment starts at a reduced angle of attack due to the negative pitch rate, which increases the effective angle of attack.
@@
The complex and transient nature of dynamic stall presents a significant obstacle in modeling, simulation, and design of aerodynamic bodies operating in dynamic stall conditions (Gardner et al., 2023; Baldan et al., 2024). Research efforts primarily focus on developing robust high-fidelity Computational Fluid Dynamics (CFD) tools, and low-fidelity reduced order models (ROMs) for implementing flow control strategies. Indeed, a deeper understanding of dynamic stall and its impact allows for improving the design and control of aerodynamic bodies (Gardner, Wolf, and Raffel, 2019; Hariharan, Egolf, and Sankar, 2014). This will ultimately lead to e.g. safer and more efficient flight, improved wind energy generation, and robust performance across diverse operating conditions.
@@
A large number of contributions in the open literature rely on the Unsteady Reynolds-Averaged Navier-Stokes (URANS) models and hybrid RANS/LES techniques to study dynamic stall (Kim and Xie, 2016; Karbasian and Kim, 2016; Khalifa, Rezaei, and Taha, 2021, 2023; Geng et al., 2018; Zanotti and Gibertini, 2013; Zanotti et al., 2014a, b; Avanzi et al., 2022, 2021; Gardner and Richter, 2020; Baldan and Guardone, 2024a, b). Most of these works conclude that the discrepancies between experimental data and numerical simulations rely on the adopted turbulence model to close the RANS system. A vast number of turbulence models have been adopted ranging from one equation Spalart-Allmaras formulation (Spalart and Allmaras, 1992) to full Reynolds stress model (RSM) (Launder, Reece, and Rodi, 1975), and including the well-known $k-\epsilon$ (Launder and Spalding, 1974) and $k-\omega$ SST (Menter, 1994). Also, transition models are included, for instance, the recent $k-\omega$ SST with intermittency equation (Menter et al., 2004). Regarding hybrid strategies, common choices are detached eddy simulations (DES) (Spalart et al., 1997) and their improved versions (DDES, IDDES). More recent approaches are scale adaptive simulations (SAS) (Menter and Egorov, 2010) and stress-blended eddy simulations (SBES) (Menter, 2018).
Large Eddy Simulations (LES) have been applied to investigate dynamic stall. In particular, several aspects have been numerically investigated for finite wings ranging from aspect ratio (Hammer, Garmann, and Visbal, 2022) to sweep angle (Hammer, Garmann, and Visbal, 2023) and compressibility effects (Benton and Visbal, 2020). Diverse studies concern nominal infinite wings obtained by imposing periodic boundary conditions in the spanwise direction. According to the best practice DES, presented in Shur et al. (1999), a spanwise size of at least one chord is required to reproduce the DSV accurately. Indeed, it is well known from experimental investigations that the characteristic dimension of the DSV is on the order of the airfoil chord length when incurring in deep dynamic stall (McCroskey, 1981). Despite the considerations mentioned, a significant portion of LES literature employs extruded airfoil geometries with spanwise lengths ranging from only 5% to 20% of the chord size. In Guillaud, Balarac, and Goncalvès (2018), a pitching NACA0012 airfoil at Re $=2\cdot 10^{4}$ is investigated using a numerical model with spanwise length of $0.5c$. In particular, the analysis compared the reduced frequency of the aerodynamic coefficients drop delay to the static case. Studies on dynamic stall phenomenology and control have been proposed by Benton and Visbal (2019a, b); Visbal and Benton (2018) and leverage high-frequency control to mitigate the intense burst of the aerodynamic loads due to the Leading Edge Vortex (LEV) detachment. These investigations also focused on the effect of the Reynolds number, which is on the order of $10^{6}$, on the laminar separation bubble (LSB). However, the wing spanwise size in all contributions is limited to $0.05c$, $0.1c$ and $0.2c$. Recently, Lee and Gross (2024) proposed to replicate LES results presented in the literature by reproducing a ramp-up motion over a straight wing with spanwise size equal to $0.1c$ and a sinusoidal pitching test case using a different geometry with span-to-chord ratio of 0.02. The same authors analyzed, in Lee, Chanez, and Gross (2024), the effects of Reynolds and Mach numbers on dynamic stall onset for harmonically pitching airfoils with LES but still limiting the span size of the model to $0.1c$. Lee and Gross (2024) work agrees with the reference data of Gupta and Ansell (2019), while Lee, Chanez, and Gross (2024) results show significant discrepancies in the downstroke phase compared to Benton and Visbal (2019a). Visbal and Garmann (2018), at the beginning of the work, focused on the span effect starting from $0.1c$ and up to $1.6c$ in a ramp-up motion. They analyzed the discrepancies in the loads and the span-averaged pressure distribution over the wing. They concluded that there is a dependency on the span size after the DSV interaction with the Trailing Edge Vortex (TEV). Despite the initial considerations on the model span extension, the core investigation focused on the LSB and DSV generation, and all the following displayed results refer to the $0.1c$ case. Another contribution by the same authors, Visbal and Garmann (2019), concerns the numerical simulations of typical experimental setups. Namely, the aspect ratio of the wing section, together with end walls and end tips, is considered. They provide a quantification of the discrepancy magnitude among all the configurations, but they do not report any comparison with actual experimental data. Less recent contributions tried to match experiments with LES. Patil et al. (2018) delves into analyzing an H-Darrieus rotor to match wind tunnel data over an entire revolution. They captured the attached flow behavior but not the stalled region. Their analysis limited the span size to $0.05c$. The contributions of Badoe, Xie, and Sandham (2019); Kasibhotla and Tafti (2014) investigate pitching NACA0012 airfoils at $2\cdot 10^{4}$ and $10^{5}$ Reynolds number using geometries with spanwise lengths of $0.25c$ and $0.18c$, respectively.
Despite recent efforts, resorting to LES or DES models do not reduce the discrepancy observed between numerical simulations and experimental data.
@@
The present work presents a numerical investigation of a NACA0012 airfoil undergoing a full sinusoidal pitching motion at a Reynolds number of 135 000 and a reduced frequency of $k=0.1$, using a wall-resolved large eddy simulation approach. The goal is to analyze the influence of spanwise length on the aerodynamic behavior during the upstroke and downstroke phases during a full sinusoidal motion. Span sizes ranging from $0.2c$ to $1.2c$ are examined to assess the dependence of the dynamic stall evolution on the span size. Numerical results are compared with the experimental data proposed by Lee and Gerontakos (2004). A comparison is presented with current state-of-the-art RANS and hybrid RANS/LES models. This comparative analysis aims to elucidate whether the observed discrepancies between the numerical results and the experimental data stem from the turbulence model employed or from other factors, such as model roughness.
@@ 1 Numerical setup
Figure 2: Details of the three-dimensional mesh of the wing section with span-to chord ratio of 1.2.
The present numerical simulations replicate the experimental conditions presented in Lee and Gerontakos (2004). A NACA0012 airfoil underlying a sinusoidal pitching motion defined as:
@@ 1.1 Grid generation
Table 1: 2D O-grid specifications. The size of the elements is made dimensionless by the chord size $c$.
@@
$\mathbf{N_{\xi}}$ $\mathbf{N_{\eta}}$ Nodes Quads $\mathbf{\Delta\xi_{\text{LE}}}$ $\mathbf{\Delta\xi_{\text{TE}}}$ $\mathbf{\overline{\Delta\xi}_{\text{suction}}}$ $\mathbf{\overline{\Delta\xi}_{\text{pressure}}}$
934 334 311 956 311 622 $6.67\cdot 10^{-4}$ $4.35\cdot 10^{-4}$ $1.88\cdot 10^{-3}$ $2.69\cdot 10^{-3}$
Table 2: 3D grid specifications. Size of elements is reported as a function of the chord size $c$.
Span size $\mathbf{N_{z}}$ Nodes Hexahedra $\mathbf{\Delta z}$
$0.2c$ 201   62.7M   62.1M $1.00\cdot 10^{-3}$
$0.4c$ 401 125.1M 124.3M $1.00\cdot 10^{-3}$
$0.8c$ 801 249.9M 248.6M $1.00\cdot 10^{-3}$
$1.2c$ 1201 374.7M 372.8M $1.00\cdot 10^{-3}$
@@
The CFD mesh is generated using Fidelity Pointwise v2023.2 Cadence (2023). To avoid the non-physical flow patterns caused by the typical C-grid strong anisotropic elements, especially near the trailing edge Baldan and Guardone (2024b), an O-grid is adopted in this research. Table 1 reports the details of the two-dimensional grid, where $\xi$ is the curvilinear coordinate along the profile and $\eta$ is the direction normal to the airfoil surface.
@@
The O-grid is obtained through a hyperbolic extrusion in the normal direction starting from the profile surface. The wall-normal spacing follows a geometric progression with a growth rate of 1.03. The standard values required for Wall-Resolved Large Eddy Simulation (WRLES), typically demanding resolution values Piomelli and Balaras (2002); Georgiadis, Rizzetta, and Fureby (2010) of $\xi^{+}<40$, $\eta^{+}<5$, and $z^{+}<20$, are satisfied at the wall for all grids, even for the most demanding flow condition when the acceleration, due to the high angle of attack, increases the velocity at more than twice the free-stream value. Thus, the first layer is positioned at $10^{-4}\ c$ over the surface. Another critical aspect of the mesh to capture dynamic stall phenomena is the element size at the leading and the trailing edges. The leading edge spacing is extremely important since it covers the region where the laminar separation bubble resides during the pitch-up phase, and subsequently, the dynamic stall vortex is generated. The same consideration can also be applied to the trailing edge since the stall also starts from the back of the profile. Also, for this reason, the blunt trailing edge is joined with a circle that avoids the imposition of the separation point and, at the same time, increases the orthogonality of the grid, allowing faster convergence. The element size at the nose of the airfoil is $\Delta\xi_{\text{leading edge}}=6.67\cdot 10^{-4}\ c$ while at the trailing edge is $\Delta\xi_{\text{trailing edge}}=4.35\cdot 10^{-4}\ c$. To reduce the number of points of the meshes, the pressure side of the airfoil, which is not directly involved in the DSV advection, is discretized with a lower resolution. Specifically, the suction side has 550 nodes, and the pressure side is limited to 385. This results in an average spacing on the upper surface of $\overline{\Delta\xi}_{\text{suction}}=1.88\cdot 10^{-3}\ c$ and on the lower surface of $\overline{\Delta\xi}_{\text{pressure}}=2.69\cdot 10^{-3}\ c$. In all the meshes, the farfield is placed at $50c$ from the airfoil to reduce the boundary influences.
@@
The three-dimensional grid is obtained through a uniform extrusion of the two-dimensional O-grid in the $z$ direction. The number of points is retrieved respecting the LES requirements (Choi and Moin, 2012). For the smallest model, extending for $0.2c$ in span, the number of points equals 201. When moving to larger extrusions, the element size is kept the same, resulting in 401, 801, and 1201 nodes for meshes with span sizes of $0.4c$, $0.8c$, and $1.2c$, respectively. Table 2 reports the details of the three-dimensional meshes. Figure 2 illustrates the three-dimensional grid with $1.2c$ span size.